# A new bump in the night: Evidence of a new feature in the binary black hole mass distribution at $70~{}M_{\odot}$ from gravitational-wave observations

### v1

@@ 5 Does spin tell us anything?
From a population point of view, the distribution of $\chi_{\rm{eff}}$ can provide hints as to the origin and formation channels of BBH mergers. A positively skewed $\chi_{\rm{eff}}$ distribution is consistent with the majority of mergers having positively aligned spins, hinting at formation in the field (Gerosa et al., 2018; Zaldarriaga et al., 2018; Belczynski et al., 2020), whereas a distribution which is symmetric about zero, hints at dynamically formed binaries due to the spin tilts being isotropically distributed at binary assembly (Rodriguez et al., 2015; Antonini & Rasio, 2016; Gerosa & Berti, 2017).
Thus to probe the origin of the sub-populations discussed in Sec. 3 for the PLPLB model, we consider extended models $p(m,\chi_{\rm{eff}}|\Lambda)$ which fit the population distribution of $\chi_{\rm{eff}}$ for each component in our mixture models. We model the $\chi_{\rm{eff}}$ distribution as a truncated Gaussian distribution in the range $\chi_{\rm{eff}}\in[-1,1]$ for each sub-population, that is, we have three distinct $\chi_{\rm{eff}}$ distributions with population parameters $\{\mu_{\chi,1},\sigma_{\chi,1},\mu_{\chi,2},\sigma_{\chi,2},\mu_{\chi,\rm{g}},\sigma_{\chi,\rm{g}}\}$ describing the location and width of the $\chi_{\rm{eff}}$ distributions for the first and second power law, as well as the Gaussian bump, respectively.
In Fig. 6, we show the posterior distribution on the $\chi_{\rm{eff}}$ population distribution parameters and in Fig. 7, we show the reconstructed primary mass and $\chi_{\rm{eff}}$ population distributions for each of the components for the PLPLB model. From our results, we find that $\mu_{\chi,1}>0$ at $>99.8$% confidence for the first power law (PL1) component, whereas for the second power law PL2 we find $\mu_{\chi,2}$ and $\mu_{\chi,\rm{g}}$ to be constrained to the within the ranges $[-0.06,0.26]$ and $[-0.92,0.41]$ respectively at 90% confidence. We find this population to be broad but skewed towards negative $\chi_{\rm{eff}}$. However, more data in this region of parameter space is needed to make definitive conclusions regarding the origin of BHs in this sub-population.
@@
Figure 7: Reconstructed $m_{1}$ and $\chi_{\rm{eff}}$ population distributions for each of the components in the PLPLB model.
6 Astrophysical Implications
7 Low-mass $\sim 10-20~{}M_{\odot}$ population
@@
We start by discussing the first power law, PL1, which appears to be consistent with the expectations for isolated binaries (e.g. van Son et al. 2022), and in agreement with findings from previous works searching for isolated binaries from this sub-population using GWTC-3 (Godfrey et al., 2024; Li et al., 2024; Ray et al., 2024). This population is likely to drop more steeply with mass than what observed in (Abbott et al., 2021b) for the broader power law, with a maximum mass buried under PL2. Its effective spin distribution is likely (at 90% CI) centered at values larger than zero, therefore supporting marginally aligned (with the binary angular momentum) spins and/or slowly spinning black holes. Aligned spins are typically predicted within the isolated binary formation channel which is in agreement with our observations. On this other hand, assuming this origin, the relatively low values may imply efficient angular momentum transport between the envelope and the core of the progenitor star leading to black holes that are slowly spinning (Bavera et al., 2020; Belczynski et al., 2020), and supernova kicks introducing misalignment between spins and binary angular momentum (Kalogera, 1996). Note that other formation channels are also expected to produce BBHs within this population, so we cannot assume that the entirety of this population (comprising $\sim 77\%$ of the BBH population according to the PLPLB model) originates in the field.
7.1 Intermediate-mass $\sim 20-40~{}M_{\odot}$ population
@@
Concerning the PL2 population, we are interpreting the $\sim 30~{}M_{\odot}$ peak as the onset of a second BH population, rather than a special feature on its own. The PLPLB model favors a maximum mass at $\sim 45-50~{}M_{\odot}$, while the PLB reaches $\sim 100~{}M_{\odot}$. Of course in the latter case, the model needs to also fit for the high mass mergers with component masses at $\sim 60~{}M_{\odot}$, so the second PL needs a higher maximum mass. Interestingly, the maximum mass for the PLPLB is $m_{\rm{max},2}=58^{+30}_{-14}~{}M_{\odot}$, and is consistent with the expectation for the onset of the pair–instability supernova (PISN) mass gap (Farmer et al., 2019), in which case it is possible that the PL2 population also originates from isolated binaries, or is made of first generation dynamically-formed binaries.
@@
For PL2 from PLPLB, we find a smoothing kernel of $\sim 20~{}M_{\odot}$, revealing a less narrow feature than in the PLB model (found at $32^{+1.9}_{-2.0}~{}M_{\odot}$ with width $3^{+4.0}_{-1.9}~{}M_{\odot}$ in the PLB model) which may be more easily explained by popular formation channels, which typically do not predict such a narrow feature at $\sim 30-35~{}M_{\odot}$.
@@
The spin distribution is relatively inconclusive with respect to the origin of PL2. Although consistent with being centered at zero, it is still marginally skewed towards aligned spins - possibly because it originates from dynamical channels but is contaminated by objects which belong to the lower mass population, or because this sub-population also originates in the field. The larger width of the distribution compared to PL1 is compelling for a dynamical origin. Biscoveanu et al. (2022); Heinzel et al. (2024b) finds a correlation between the width of the effective spin distribution and redshift. We argue that this may be related to what we observe here. If we allow for the different sub-populations to vary with redshift independently of each other, we find that at higher redshift the lower mass population rate decreases (as also found by Rinaldi et al. 2024) while the PL2 rate increases. While such redshift evolution may not be physical due to the limited detector horizon especially at lower masses (Heinzel et al., 2024a), this finding may reconcile our and previous works’ observations that the second sub-population drives the spreading of the spin width at higher redshift. However, note that uncertainties on spins for higher mass systems are typically larger than those from PL1, so it is challenging to draw definitive conclusions. Perhaps the most compelling evidence that PL2 may be composed of isolated binaries or dynamically formed first generation binaries is the fact that the PLPLB model PL2 has a maximum mass consistent with PISN mass gap. An involve study exploring the redshift and $\chi_{\rm{eff}}$ dependence on the mass distribution is ongoing.
@@
We note that Godfrey et al. (2024) and (Li et al., 2024) also find evidence for 2 distinct sub-populations. Specifically, Godfrey et al. (2024) finds that the higher mass population which we identify here with PL2 shares characteristics with the likely isolated binary population peaking at $\sim 10~{}M_{\odot}$. Given all of the arguments and previous works outlined, we argue that PL1 is composed of isolated binaries, while PL2 may mostly arise from 1g+1g binaries from dynamical formation.
7.2 High-mass $\sim 60-80~{}M_{\odot}$ population
@@
Although it may be tempting to identify the feature at $\sim 70~{}M_{\odot}$ as the pulsational PISN (PPISN) build-up (Talbot & Thrane, 2018), the bump is located at a mass too high to be at the lower edge of the PISN mass gap, which is expected to be around $45~{}M_{\odot}$ and likely below $56~{}M_{\odot}$ considering reasonable variations in metallicity and uncertainties in the CO reaction rates (Farmer et al., 2019). For the high mass bump, we find a broad distribution on $\chi_{\rm{eff}}$. Although with large uncertainties given the current statistics, this population supports the presence of effective spins $<-0.3$, which are expected to exist for hierarchical mergers (Baibhav et al., 2020; Fishbach et al., 2022), but are hardly explained by isolated binary formation. Given these arguments, we argue that a plausible explanation for this feature is from hierarchical mergers of BHs from the $\sim 30-35~{}M_{\odot}$ bump. Because we find that the fraction of BBHs in the bump is 1-4% at 90% confidence, assuming these are from 2g BHs in dense star clusters, it is likely that these BHs are born spinning ($\chi_{\rm birth}>0$), but not enough to wash out this feature completely ($\chi_{\rm birth}<0.5$), according to the predictions of Rodriguez et al. 2019. A high mass distribution similar to what found here is also predicted by Kremer et al. (2020), with the bump above $\sim 60~{}M_{\odot}$ originating from 2g hierarchical mergers and BHs born from stellar collision.
8 Conclusions
@@
We also find that adding a Gaussian component above $40~{}M_{\odot}$ robustly finds a sub-population located around $65-70~{}M_{\odot}$. It is possible that by allowing a second power law to exist with a steeper slope than the first PL, one enables the detection of a sub-population that was otherwise buried under a single, flatter power law. This sub-population may arise from second generation hierarchical mergers, forming from black holes in the secondary peak of the distribution at $\sim 35~{}M_{\odot}$. Our simulations show that a similar feature cannot be recovered by chance coincidence when the underlying population follows a PLB. However, more data is required to confidently claim that the PLPLB model is preferred over other models. We expect that the increasing number of BBH events detected through O4 and beyond will provide the observations required to determine whether the $70~{}M_{\odot}$ sub-population is actually present in the astrophysical BH mass distribution. If confirmed, such sub-population would a smoking gun about the dynamical origin of high-mass LVK BBHs, and may represent a new valuable feature to improve cosmological parameters constraints through spectral standard siren measurements (Farr et al., 2019; Ezquiaga & Holz, 2022).
9 Acknowledgements

### v2

@@ V Does spin tell us anything?
From a population point of view, the distribution of $\chi_{\rm{eff}}$ can provide hints as to the origin and formation channels of BBH mergers. A positively skewed $\chi_{\rm{eff}}$ distribution is consistent with the majority of mergers having positively aligned spins, hinting at formation in the field (Gerosa et al., 2018; Zaldarriaga et al., 2018; Belczynski et al., 2020), whereas a distribution which is symmetric about zero, hints at dynamically formed binaries due to the spin tilts being isotropically distributed at binary assembly (Rodriguez et al., 2015; Antonini and Rasio, 2016; Gerosa and Berti, 2017).
Thus to probe the origin of the sub-populations discussed in Sec. III for the PLPLB model, we consider extended models $p(m,\chi_{\rm{eff}}|\Lambda)$ which fit the population distribution of $\chi_{\rm{eff}}$ for each component in our mixture models. We model the $\chi_{\rm{eff}}$ distribution as a truncated Gaussian distribution in the range $\chi_{\rm{eff}}\in[-1,1]$ for each sub-population, that is, we have three distinct $\chi_{\rm{eff}}$ distributions with population parameters $\{\mu_{\chi,1},\sigma_{\chi,1},\mu_{\chi,2},\sigma_{\chi,2},\mu_{\chi,\rm{g}},\sigma_{\chi,\rm{g}}\}$ describing the location and width of the $\chi_{\rm{eff}}$ distributions for the first and second power law, as well as the Gaussian bump, respectively.
In Fig. 7, we show the posterior distribution on the $\chi_{\rm{eff}}$ population distribution parameters. In Fig. 8, we show the reconstructed primary and secondary mass population distributions for each of the components for the PLPLB model. Similarly, in Fig. 9 we show the reconstructed $\chi_{\rm{eff}}$ population distribution for each of the components for the PLPLB model. From our results, we find that $\mu_{\chi,1}>0$ at $>99.8$% confidence for the first power law (PL1) component and to be constrained within the range $[0.02,0.08]$ at 90% confidence. Whereas for the second power law (PL2) and Gaussian Bump we find $\mu_{\chi,2}$ and $\mu_{\chi,\rm{g}}$ to be constrained to the within the ranges $[-0.08,0.15]$ and $[-0.87,0.52]$ respectively at 90% confidence. We find this population to be broad but skewed towards negative $\chi_{\rm{eff}}$. However, more data in this region of parameter space is needed to make definitive conclusions regarding the origin of BHs in this sub-population.
@@
Figure 8: Reconstructed $m_{1}$ and $m_{2}$ population distributions for each of the components in the PLPLB model.
Figure 9: Reconstructed $\chi_{\rm{eff}}$ population distributions for each of the components in the PLPLB model. We also show the $\text{CDF}(\chi_{\rm{eff}})$ to more easily demonstrate the symmetry (about $\chi_{\rm{eff}}=0$) for each sub-population.
VI Astrophysical Implications
VI.1 Low-mass $\sim 10-20~{}M_{\odot}$ population
@@
We start by discussing the first power law, PL1, which appears to be consistent with the expectations for isolated binaries (van Son et al., 2022), and in agreement with findings from previous works searching for isolated binaries from this sub-population using GWTC-3 (Godfrey et al., 2024; Li et al., 2024; Ray et al., 2024). This population is likely to drop more steeply with mass than what observed in (Abbott et al., 2021b) for the broader power law, with a maximum mass buried under PL2. Its effective spin distribution is likely (at 90% CI) centered at values larger than zero, therefore supporting marginally aligned (with the binary angular momentum) spins and/or slowly spinning black holes. Aligned spins are typically predicted within the isolated binary formation channel which is in agreement with our observations. On this other hand, assuming this origin, the relatively low values may imply efficient angular momentum transport between the envelope and the core of the progenitor star leading to black holes that are slowly spinning (Bavera et al., 2020; Belczynski et al., 2020), and supernova kicks introducing misalignment between spins and binary angular momentum (Kalogera, 1996). Note that other formation channels are also expected to produce BBHs within this population, so we cannot assume that the entirety of this population (comprising $\sim 76\%$ of the BBH population according to the PLPLB model) originates in the field.
VI.2 Intermediate-mass $\sim 20-40~{}M_{\odot}$ population
@@
Concerning the PL2 population, we are interpreting the $\sim 30~{}M_{\odot}$ peak as the onset of a second BH population, rather than a special feature on its own. The PLPLB model favors a maximum mass at $\sim 45-50~{}M_{\odot}$, while the PLB reaches $\sim 100~{}M_{\odot}$. Of course in the latter case, the model needs to also fit for the high mass mergers with component masses at $\sim 60~{}M_{\odot}$, so the second PL needs a higher maximum mass. Interestingly, the maximum mass for the PLPLB is $m_{\rm{max},2}=58^{+32}_{-14}~{}M_{\odot}$, and is consistent with the expectation for the onset of the pair–instability supernova (PISN) mass gap (Farmer et al., 2019; Farag et al., 2022), in which case it is possible that the PL2 population also originates from isolated binaries, or is made of first generation dynamically-formed binaries.
@@
For PL2 from PLPLB, we find a smoothing kernel of $\sim 40~{}M_{\odot}$, revealing a less narrow feature than in the PLB model (found at $32.0^{+1.9}_{-2.0}~{}M_{\odot}$ with width $3.1^{+3.9}_{-1.8}~{}M_{\odot}$ in the PLB model) which may be more easily explained by popular formation channels, which typically do not predict such a narrow feature at $\sim 30-35~{}M_{\odot}$.
@@
The spin distribution is relatively inconclusive with respect to the origin of PL2. Although consistent with being centered at zero, it is still marginally skewed towards aligned spins - possibly because it originates from dynamical channels but is contaminated by objects which belong to the lower mass population, or because this sub-population also originates in the field. The larger width of the distribution compared to PL1 is compelling for a dynamical origin. Recent work (Biscoveanu et al., 2022; Heinzel et al., 2024a) finds a correlation between the width of the effective spin distribution and redshift. We argue that this may be related to what we observe here. If we allow for the different sub-populations to vary with redshift independently of each other, we find that at higher redshift the lower mass population rate decreases as also found by (Rinaldi et al., 2024), while the PL2 rate increases. While such redshift evolution may not be physical due to the limited detector horizon especially at lower masses (Heinzel et al., 2024b), this finding may reconcile our and previous works’ observations that the second sub-population drives the spreading of the spin width at higher redshift. However, note that uncertainties on spins for higher mass systems are typically larger than those from PL1, so it is challenging to draw definitive conclusions. Perhaps the most compelling evidence that PL2 may be composed of isolated binaries or dynamically formed first generation binaries is the fact that the PLPLB model PL2 has a maximum mass consistent with PISN mass gap. An involved study exploring the redshift and $\chi_{\rm{eff}}$ dependence on the mass distribution is ongoing.
@@
We note that (Godfrey et al., 2024) and (Li et al., 2024) also find evidence for 2 distinct sub-populations. Specifically, (Godfrey et al., 2024) finds that the higher mass population which we identify here with PL2 shares characteristics with the likely isolated binary population peaking at $\sim 10~{}M_{\odot}$. Given all of the arguments and previous works outlined, we argue that PL1 is composed of isolated binaries, while PL2 may mostly arise from 1g+1g binaries from dynamical formation.
VI.3 High-mass $\sim 60-80~{}M_{\odot}$ population
@@
Although it may be tempting to identify the feature at $\sim 70~{}M_{\odot}$ as the pulsational PISN (PPISN) build-up (Talbot and Thrane, 2018), the bump is located at a mass too high to be at the lower edge of the PISN mass gap, which is expected to be around $45~{}M_{\odot}$ and likely below $56~{}M_{\odot}$ considering reasonable variations in metallicity and uncertainties in the CO reaction rates (Farmer et al., 2019). For the high mass bump, we find a broad distribution on $\chi_{\rm{eff}}$. Although with large uncertainties given the current statistics, this population supports the presence of effective spins $<-0.3$, which are expected to exist for hierarchical mergers (Baibhav et al., 2020; Fishbach et al., 2022), but are hardly explained by isolated binary formation. Given these arguments, we argue that a plausible explanation for this feature is from hierarchical mergers of BHs from the $\sim 30-35~{}M_{\odot}$ bump. Because we find that the fraction of BBHs in the bump is 1-5% at 90% confidence, assuming these are from 2g BHs in dense star clusters, it is likely that these BHs are born spinning ($\chi_{\rm birth}>0$), but not enough to wash out this feature completely ($\chi_{\rm birth}<0.5$), according to the predictions of  (Rodriguez et al., 2019). A high mass distribution similar to what found here is also predicted by (Kremer et al., 2020), with the bump above $\sim 60~{}M_{\odot}$ originating from 2g hierarchical mergers and BHs born from stellar collision.
VII Conclusions
@@
We also find that adding a Gaussian component above $40~{}M_{\odot}$ robustly finds a sub-population located around $65-70~{}M_{\odot}$. It is possible that by allowing a second power law to exist with a steeper slope than the first PL, one enables the detection of a sub-population that was otherwise buried under a single, flatter power law. This sub-population may arise from second generation hierarchical mergers, forming from black holes in the secondary peak of the distribution at $\sim 35~{}M_{\odot}$. Our simulations show that a similar feature cannot be recovered by chance coincidence when the underlying population follows a PLB. However, more data is required to confidently claim that the PLPLB model is preferred over other models. We expect that the increasing number of BBH events detected through O4 and beyond will provide the observations required to determine whether the $70~{}M_{\odot}$ sub-population is actually present in the astrophysical BH mass distribution. If confirmed, such sub-population would a smoking gun about the dynamical origin of high-mass LVK BBHs, and may represent a new valuable feature to improve cosmological parameters constraints through spectral standard siren measurements (Farr et al., 2019; Ezquiaga and Holz, 2022).
VIII Acknowledgements